OpenStudy (anonymous):

Solve the triangle. A = 46°, a = 33, b = 26 anyone out there want to help a lost soul?

OpenStudy (anonymous):

did you draw one?

OpenStudy (anonymous):

the first step is to draw a reasonable looking triangle

OpenStudy (anonymous):

 \[\frac{ 33 }{ \sin46 } =\frac{ 26 }{ sinB}\] \[SinB =\frac{ \sin46 * 26 }{ 33 } = 0.56\] B= \[\sin^{-1} 0.56 =34. 52° angle for C = 180°-46°-34.5° = 99.5°.

OpenStudy (anonymous):



OpenStudy (anonymous):

@lap. may arrive at the correct answer, but i have to point out that the triangle is not quite right

OpenStudy (anonymous):

so \[\frac{ 33 }{ \sin46 } =\frac{ c }{ \sin 99.5 }\] \[c = \frac{ 33 *\sin99.5 }{ \sin 46 } =45.25\]

OpenStudy (anonymous):

see for this one I solved for angle B by setting up a triangle but I have no idea how to solve for side c to get angle C

OpenStudy (anonymous):

wait how'd you get angle C?

OpenStudy (anonymous):

see previous post angle for C = 99.5° and you have to use The Law of Sines : http://www.mathsisfun.com/algebra/trig-sine-law.html

OpenStudy (anonymous):

thanks!! i understand now.

OpenStudy (anonymous):

you are welcome :)